current dividend is $2; current stock price is $30.00; future growth of dividend is 6%. what is this firm's cost of equity using the DCF approach?
Answer #1

This question requires application of constant growth dividend discount model according to which: Po - Divi T-9 Po = Price of

Div1 = $2 * (1 + 6%) = $2.12

30 = 2.12 T-006

r - 0.06 = 0.0707

r = 13.07%